# A Comprehensive Simulation Framework for CXL Disaggregated Memory

### v3

@@ IV-B4 Real-World Application Test
The benchmarks above show that the CXL-FPGA, operating at a lower frequency, cannot accurately replicate the actual performance of real CXL memories. Due to limitations on space and simulation time, future real-world applications will only use the CXL-ASIC devices. With the above characterization results, we can see that the performance of CXL-FPGA is much poorer than the real CXL-ASIC device. Due to the simulation overhead as well as page limitations, we will limit our real-world application experiments to CXL-ASIC only. To better understand the system performance of CXL-based disaggregated memory, we selected a representative in-memory database named Redis [39] to evaluate the system throughput in diverse application scenarios. The performance of Redis is evaluated using YCSB, a popular and well-known NoSQL and SQL database benchmark suite with six predefined workloads [40].
@@
Fig. 13 shows the performance metric Queries Per Second (QPS) of Redis under two workloads with different memory allocation strategies. For both workloads, using only DDR-L memory results in the highest QPS, while 100% CXL or DDR-R memory leads to smallest QPS. This is because mixing poor-performance (prominently in latency as observed previously) memory into DDR-L pulls down the overall performance of Redis. In addition, the QPS achieved with a certain combination of DDR-L and DDR-R is slightly higher than that achieved with the same combination of DDR-L and CXL memories. Redis is a memory-latency-sensitive application which operates with -level precision[25, 45, 46]. For such applications, even a minor allocation of their working set to high-latency memory can lead to significant performance degradation. Therefore, latency-bound applications such as Redis, DDR-R may deliver superior performance compared to CXL-ASIC due to its moderate latency characteristics.
@@ IV-C1 Memory Capacity Expansion
Fig. 14 shows the test results of system performance in normalized QPS. When inserting key-value pairs of <16,200>, the QPS remains relatively consistent across various memory configurations. This is because the total data volume inserted into Viper did not exceed the maximum capacity of DDR-L during the execution of the four operations. However, when the size of the inserted key-value pairs increased to <100,900>, the system’s QPS with DDR-L alone dropped significantly. The decline is attributed to the total volume of inserted data surpassing the capacity threshold that DDR-L can accommodate. It cannot completely accommodate the data intended to be inserted into Viper, necessitating frequent swapping of pages in and out using swap space by default. It can be seen that expanding memory with CXL memory can enhance system throughput significantly with at most 25 times or 23 times improvement compared to DDR-L alone. This suggests that CXL memory can be an effective means of expanding system memory on top of existing DDR memory when needed, as it does not occupy DIMM slots and simply relies on PCIe 5.0/6.0 x16 interface.
@@ IV-C2 Memory Bandwidth Expansion
To evaluate the benefits of CXL memory to bandwidth-sensitive applications, we performed inference tasks using Facebook’s Deep Learning Recommendation Model (DLRM) configured similar to the prior work MERCI [48]. We conducted tests with various memory mixing configurations, covering DDR-L, DDR-R, and CXL HDM with an interleaved memory allocation strategy.
Fig. 15(a) presents the test results of DLRM inference in various memory mixing scenarios. One can see that the system performance improves 29% and 16% separately when using a 50%:50% mix of DDR-L and DDR-R/CXL-HDM in comparison to the 100% DDR-L case. This is because interleaved memory can effectively expand the bandwidth of the existing DDR memory, thereby increasing the overall system’s QPS. When using a 50%:50% mix of DDR-L and CXL HDM, the QPS of the system is lower than the QPS achieved with a 50%:50% mix of DDR-L and DDR-R. This can be explained by the fact that the emulated CXL memory using remote DDR5 offers a relatively higher performance considering the combined effects of both latency and bandwidth when compared to the CXL memory.
The simulation results on CXL-DMSim suggest a similar phenomenon, despite the QPS is slightly lower than its counterpart on real hardware. The variation can be attributed to the fact that the X86O3CPU model we used on CXL-DMSim cannot absolutely match today’s advanced Xeon CPU. We further analyzed the root causes of the phenomenon using detailed output statistics of the simulator. In Fig.15(b), the metric dram.avgQLat represents the average queue latency per DRAM burst for the DDR memory, whereas dram.avgMemAccLat is the average memory access latency per DRAM burst. l3cache.overallAvgMissLatency measures the average cache miss latency. The statistics reveal that hybrid DDR and CXL memories deliver the lowest latency across all metrics, due to reduced queuing delays and access contention. Note that in the CXL-ASIC exclusive configuration, there is no DDR DRAM traffic; thus dram.avgMemAccLat and dram.avgQLat are both 0.
@@ IV-D Congestion Analysis Using CXL-DMSim
During our experiments, we observed an interesting phenomenon that after a certain point the QPS for MERCI drops as the number of cores continuously increases. Using performance counter monitor (PCM) [49], we found that the average L3 cache miss latency of the 48-core system is , whereas the data is for the 12 cores. We intuitively speculate that under high-concurrent memory access pressure in the 48-core system, the CXL link may experience congestion, resulting in performance degradation. Fortunately, CXL-DMSim’s high observability allows us to dive into the system to further investigate what is happening and its root causes.
Table IV shows the QPS and key statistics from CXL-DMSim under different configurations, where stats.1 shows that the aggregate QPS of the 12-core system is twice that of the 48-core system. Stats.2-7 present various latency parameters for a single core. Stats.2 and 3 reveal that in the 48-core system, the average core load-to-use latency and its standard deviation are significantly higher than those in the 12-core system. Stat.4 shows the proportion of load-to-use latency within 0-9 cycles, while stats.5 and stats.6 reflect the minimum and maximum load-to-use latency, respectively. These three metrics suggest that although 94% of the loads in the 48-core system can respond quickly (within 0-9 cycles), a small fraction of the loads exhibit long-tail latency, leading to a larger standard deviation in the load-to-use latency distribution. Stats.7 records core stall events caused by a full load/store queue (LSQ). The data shows that congestion from a small number of load requests in the 48-core system leads to more frequent core stalls, resulting in a lower QPS. In contrast, while only 63% of the loads in the 12-core system can respond quickly, the smaller variance in load-to-use latency suggests a more pipelined system, which results in a higher QPS. We further localized the congestion point on the CXL path. Stats.9 records the number of transmission retries caused by full request/response queues in the bridge module, with results showing that the retry count in the 48-core system is three times higher than the 12-core system. Stats.10 reflects the average response time of the CXL device, showing that the device response times in both systems are nearly identical. In summary, the high degree of parallel memory access pressure in the 48-core system leads to severe queuing in the bridge module, resulting in long-tail latency.
@@
| 1 | Aggregate QPS | 1.8e6 | 8.8e5 |
| 2 | core.loadToUse::mean (Cycle) | 97 | 198 |
| 3 | core.loadToUse::stdev (Cycle) | 217 | 1107 |
| 4 | core.loadToUse::0-9 (Cycle) | 63.6% | 94.8% |
| 5 | core.loadToUse::min_value (Cycle) | 2 | 2 |
| 6 | core.loadToUse::max_value (Cycle) | 2.7e3 | 2.5e4 |
| 7 | core.lsqFullEvents (Count) | 4.6e6 | 9.0e6 |
| 8 | l3.overallAvgMissLat. (Tick) | 1.3e5 | 1.2e6 |
| 9 | bridge.reqRetryCounts (Count) | 1.8e7 | 6.1e7 |
| 10 | cxl.rsp::mean (Tick) | 8.1e4 | 8.1e4 |
With the above analysis, we speculate that because the CXL link is longer than DDR links, it is more prone to congestion under high memory access pressure, thereby causing the system performance to degrade rapidly. With the help of CXL-DMSim, researchers can easily pinpoint congestion locations and quickly explore various optimization strategies, such as smarter CPU prefetch mechanisms, customized flow control strategies, and more advanced CXL memory controller.
@@ IV-E Expandability of CXL-DMSim
The CXL protocol is agnostic to the underlying memory technologies, enabling seamless support for a range of media, including DRAM, Flash, and emerging non-volatile memories. CXL-DMSim also adheres to this principle, allowing for the convenient extension of internal components.
@@
As an example, Fig. 16 shows the device model of CXL-SSD, which is a newly added component to CXL-DMSim. The red dashed blocks highlight the added modules to our CXL device model. In the memory medium module, we utilized an open-source SSD simulator called SimpleSSD [50, 51] to model the SSD backend of the CXL interface. However, two issues must be addressed. First, SSD exhibits significantly higher latency ( level) compared to DRAM ( level). Second, CXL.mem is based on memory semantics at a byte granularity, while SSD operates with I/O semantics at a page (e.g., 4KB) granularity. A simple replacement of the memory medium from DRAM to SSD would inevitably result in a dramatic performance decline. Thus, we designed a cache between the device controller and memory medium to address the aforementioned issues. Additionally, we implemented a cache prefetching and replacement module to manage cache lines. Specifically, we first considered that, due to the larger access granularity of SSDs, prefetching is more likely to result in cache pollution. Our prefetching decisions are based on the history of cache misses, with prefetching only executed when there is a high confidence level that it will yield benefits. Furthermore, given the higher access latency of SSDs, we employ the Best-Offset prefetching algorithm [52], which prioritizes timeliness in retrieving data from the backend SSD. When a memory access request reaches the CXL-SSD, it is directly returned if the SSD cache hits; otherwise, it is translated into SimpleSSD data packets for backend retrieval.
@@
We ran the Viper test on the CXL-SSD to evaluate its performance. Unlike the previous tests, we utilized the AM mode in this case to demonstrate its usage. As illustrated in Fig. 17, the test results indicate that the QPS of CXL-SSD across four operations are considerably lower than CXL-DRAM, due to the higher access latency of Flash medium. To evaluate the effects of cache on system performance, we implemented and compared different cache replacement policies including Least Recently Used (LRU) and First In First Out (FIFO). It can be seen that the QPS of LRU or FIFO CXL-SSD with an additional cache has been significantly enhanced, reaching 72%-88% of CXL-DRAM. The results of this test not only prove the effectiveness of the AM mode, but also demonstrate the good expandability of our simulator. This means that CXL-DMSim can simulate various CXL memory devices integrated with different storage medium, providing a broad exploration and design space for the research of future heterogeneous memory-pooled systems.
@@ V Discussion
| Attribute | CXL-DMSim | QEMU [17] | Mess+gem5 [26] | gem5-cxl [18] | CXLMemSim [19] | Remote NUMA [21, 22, 23] |
| CXL protocol support | Yes | Yes | No | No | No | No |
| Full-system CXL support | Yes | Yes | No | No | No | No |
| Silicon validation | Yes | No | No | No | No | N/A |
| Expandability | High | High | Low | Low | Low | Low |
| Sw/hw co-design | Yes | Yes | No | No | No | No |
| Cycle accuracy | Yes | No | Yes | Yes | No | Yes |
| Development maturity | Yes | Yes | Yes | No | No | Yes |
| Configurability | High | High | High | Low | Medium | Low |
| Simulation error | Medium | High | Medium | High | Unknown | High |
| Simulation speed | Low | Medium | Low | Low | Medium | High |
@@ V-A Simulation Overhead of CXL-DMSim
To assess the simulation overhead covering two aspects which are simulation speed and occupied host memory introduced by CXL-DMSim in comparison to the raw gem5, we conducted various experiments on our simulator and collected statistics of the execution time (HostSeconds) and physical memory occupation (HostPhyMemory) on the host machine. Fig 18(a) compares the program execution time between raw gem5, CXL-DMSimL, and CXL-DMSimA with respect to three workloads. The execution time of CXL-DMSimL is nearly identical to that of raw gem5, while the average execution time of CXL-DMSimA is 16% higher. This is because CXL-DMSimA involves more components and longer latency in the simulation process. Notably, LMBench, a benchmark focused on evaluating memory access latency, exhibits slightly higher execution times compared to STREAM and MERCI.
The comparison of HostPhyMemory overhead is shown in Fig 18(b). CXL-DMSim exhibits an average increase of only 3.5% in host memory usage, compared to raw gem5. This increase results from the additional memory allocation necessary on the host machine to model the CXL device memory space. Moreover, the specific magnitude of memory overhead is closely correlated with the amount of physical memory utilized by different applications.
@@ V-B Simulation Error of CXL-DMSim
We also analyzed the simulation error stats of CXL-DMSim for a large variety of workloads against the results on our hardware testbed. The collected error data is compared with that of other CXL simulators as depicted in Fig. 19. It can be seen that the minimum simulation error is 1.4% for MERCI while the maximum goes to Viper. The averaged simulation error is measured at 4.1% across all workloads in our experiments. CXL-DMSim exhibits the best accuracy in comparison to the two latest works Mess (6.0%) and gem5Tune (6.2%) as revealed in their respective papers [26, 53]. The native gem5 shows the highest average error of 15.0%[26].
The sources of simulation errors for CXL-DMSim can be attributed to the cumulative effects of modeling inaccuracies in the entire system. These include but not limited to the following. (1) The CPU model lags behind of the latest CPU we used on our testbed. (2) The poor modeling of disk in gem5 when storage layers are involved, as can be observed with the increased errors for Viper when some data was pushed out to disk. Note that it is always a dilemma in pursuit of simulation accuracy and speed.
@@
Table V compares CXL-DMSim and other prominent simulators such as QEMU, Mess, gem5-CXL, CXLMemSim and remote NUMA at various aspects. First, CXL-DMSim and QEMU offer comprehensive support for CXL protocol, which is critical for accurately simulating the behavior of real CXL memory systems. In contrast, the other four simulators actually do not support CXL protocol. Second, CXL-DMSim is a full-system CXL simulator that provides the most realistic interaction between the operating system and simulated hardware, enabling a range of opportunities for software-hardware co-design. In contrast, Mess, gem5-CXL and CXLMemSim cannot boot the operating system that supports CXL devices. Third, CXL-DMSim is the only one which has undergone extensive silicon validation, surpassing the other simulators in terms of accuracy and realism. CXL-DMSim also excels in expandability, cycle accuracy, development maturity, configurability, with acceptable simulation error. Nevertheless, these advantages come along at the cost of slow simulation speed, high modeling complexities for new features, and limited system scale.

### v6

@@ IV-B3 Real-World Application Test
With the above characterization results, we can see that the performance of CXL-FPGA is much poorer than the real CXL-ASIC device. Due to the simulation overhead as well as page limitations, we will limit our real-world application experiments to CXL-ASIC only. To better understand the system performance of CXL-based disaggregated memory, we selected a representative in-memory database named Redis [32] (version 5.0.5) to evaluate the system throughput in diverse application scenarios. The performance of Redis is evaluated using YCSB, a popular database benchmark suite [33]. We used predefined workloads A and B, both following a Zipfian distribution. We employed YCSB’s default key-value size, an working set, and a single thread for the experiments.
@@
With numactl, we can direct Redis to allocate memory in DDR-L, DDR-R, CXL, or interleave between them to simulate a heterogeneous memory allocation strategy. Fig. 12 shows the queries per second (QPS) of Redis under two workloads with different memory allocations. Using only DDR-L yields the highest QPS (42K QPS for Workload A and 39K QPS for Workload B), while interleaved allocations reduce QPS, and 100% CXL or DDR-R leads to the lowest QPS. Additionally, an interleaving of DDR-L and DDR-R achieves slightly higher QPS than the same combination of DDR-L and CXL memories. Redis is a memory-latency-sensitive application operating at -level precision [22, 38]. For such applications, even a minor allocation of their working set to high-latency memory can lead to significant performance degradation. DDR-R, due to its moderate latency characteristics, may exhibit performance superior to that of CXL-ASIC.
@@ IV-C1 Memory Capacity Expansion
Fig. 11 shows the test results of system performance in normalized QPS. When inserting key-value pairs of <16,200>, the QPS remains relatively consistent across various memory configurations. This is because the total data volume inserted into Viper did not exceed the maximum capacity of DDR-L during the execution of the four operations. However, when the size of the inserted key-value pairs increased to <100,900>, the system’s QPS with DDR-L alone dropped significantly. The decline is attributed to the total volume of inserted data surpassing the capacity threshold that DDR-L can accommodate. It cannot completely accommodate the data intended to be inserted into Viper, necessitating frequent swapping of pages in and out using swap space by default. It can be seen that expanding memory with CXL memory can enhance system throughput significantly with at most 25 times or 23 times improvement compared to DDR-L alone.
@@ IV-C2 Memory Bandwidth Expansion
To evaluate the benefits of CXL memory to bandwidth-sensitive applications, we performed inference tasks using Meta’s Deep Learning Recommendation Model (DLRM) configured similar to the prior work MERCI [40]. We conducted tests with various memory mixing configurations, covering DDR-L, DDR-R, and CXL HDM with an interleaved memory allocation strategy.
Fig. 13(a) presents the test results of DLRM inference in various memory mixing scenarios. One can see that the system performance improves by 39% and 60% respectively, when using a 50%:50% mix of DDR-L and DDR-R/CXL-HDM compared to the 100% DDR-L case. This is because interleaved memory can effectively utilize the bandwidth of multiple memory devices, thereby increasing the overall system QPS. When using a 50%:50% mix of DDR-L and CXL HDM, the QPS of the system is higher than the QPS achieved with a 50%:50% mix of DDR-L and DDR-R. This improvement stems from the superior bandwidth of CXL-ASIC memory over DDR-R. Notably, CXL-DMSim produced simulation results consistent with real hardware observations, demonstrating its capability to accurately model the characteristics of CXL memory devices. To gain deeper insights into the root causes of bandwidth expansion in interleaved memory, we analyzed detailed statistics collected from the simulator. In Fig. 13(b), the metric llc.AvgMissLat measures the average LLC miss latency, dram.avgMemAccLat represents the average memory access latency per DRAM burst, and dram.avgQLat quantifies the average queuing latency per DRAM burst in DDR memory. Note that in the CXL-ASIC exclusive configuration, there is no DDR traffic; thus, both dram.avgMemAccLat and dram.avgQLat are zero. The statistical results demonstrate that the interleaved DDR and CXL memory configuration consistently exhibited the lowest latency across all metrics. This improvement is primarily due to CXL memory offloading half of the memory traffic, significantly reducing queuing delays and access contention. Consequently, this configuration achieved the lowest LLC average miss latency.
@@ IV-D Congestion Analysis Using CXL-DMSim
During our experiments, we observed an interesting phenomenon in both CXL-DMSim and the hardware platform. After a certain point, the QPS for MERCI drops as the number of cores continuously increases. Using the performance counter monitor (PCM) [41] on the hardware testbed, we found that the average LLC miss latency in the 48-core system was , compared to in the 12-core system. Due to the limited observability of the real hardware, we used CXL-DMSim to dig into the root causes. We collected key statistics for both the simulated 12-core and 48-core systems, as listed in Table IV. Stats.1 shows that the aggregate QPS of the 48-core system is only 69% of that of the 12-core system. Meanwhile, Stats.2 and Stats.3 reveal that both the average per-core load-to-use latency and the average LLC miss latency are higher in the 48-core system than in the 12-core system, highlighting the increased latency along the CXL memory access path in the 48-core configuration. Stats.4 records the count of blocked events of LLC requests caused by miss status handling register (MSHR) entry exhaustion. As MSHRs are critical for managing concurrent cache miss requests, their depletion blocks subsequent CPUs from continuing to issue memory access requests, significantly degrading system performance.
on the CXL-DMSimA memory (Cfg9).
@@
| 1 | Aggregate QPS | 1.6e6 | 1.1e6 |
| 2 | core.loadToUse::mean (Cycle) | 97 | 143 |
| 3 | llc.overallAvgMissLat. (Tick) | 3.1e5 | 9.6e5 |
| 4 | llc.mshrBlocked (Count) | 0 | 1.6e7 |
| 5 | bridge.avgReqQueOccupancy | 22.6% | 43.0% |
| 6 | bridge.avgRspQueOccupancy | 48.7% | 99.8% |
| 7 | cxlCtrl.avgReqQueOccupancy | 13.6% | 13.9% |
| 8 | cxlCtrl.avgRspQueOccupancy | 66.3% | 99.0% |
| 9 | cxlMem.avgRdBufOccupancy | 14.4% | 21.9% |
| 10 | cxlMem.avgWrBufOccupancy | 37.8% | 38.2% |
To further analyze the root cause, we conducted an in-depth examination of packet behaviors along the CXL transmission path, which consists of four components: the bridge, the CXL controller, memory controller, and the memory medium. In our CXL link implementation, reliable sequential transmission and credit-based flow control [42] follow two key constraints: 1) a requester must ensure the receiver’s request queue has sufficient capacity (credits) before sending packets; 2) for requests requiring responses, the receiver must pre-allocate enough space in the response queue to prevent deadlock (e.g., deadlock occurs when subsequent requests depend on the results of preceding requests, but these responses cannot be returned due to a full response queue, thereby blocking further requests). Stats.5-10 quantify the average resource utilization of different components along the CXL transmission path. Compared to the 12-core system, we observed severe packet queuing congestion in the response queues of both the bridge and the CXL controller in the 48-core system. Meanwhile, the occupancy of the read/write buffers in the backend memory module, though slightly higher, remains below 40% of the full capacity. These observations suggest that the congestion primarily originates from the CXL controller rather than the memory module. Specifically, in the 48-core system, highly concurrent memory accesses rapidly depleted the available credits for the response queue in the CXL controller. This congestion then causes back-pressure on the bridge and LLC, ultimately exhausting LLC MSHR entries and thus blocking further CPU memory accesses, thereby degrading system QPS.
@@ IV-E Expandability of CXL-DMSim
CXL-DMSim adheres to the memory technology-agnostic principle of the CXL protocol, enabling seamless support for a rang of media such as DRAM, Flash, and emerging NVMs. As an example, Fig. 14 shows the device model of CXL-SSD, which is a newly added component to CXL-DMSim. The red dashed blocks highlight the added modules to our CXL device model. We utilized an open-source SSD simulator called SimpleSSD [43] to model the SSD backend of the CXL interface. However, two issues must be addressed. First, SSD exhibits significantly higher latency ( level) compared to DRAM ( level). Second, CXL.mem is based on memory semantics at a byte granularity, while SSD operates with I/O semantics at a page (e.g., 4KB) granularity. A simple replacement of the memory medium from DRAM to SSD would inevitably result in a dramatic performance decline. Thus, we designed a cache between the CXL controller and SSD model to address the aforementioned issues. Additionally, we implemented a cache prefetching and replacement module to manage cache lines. Specifically, we first considered that, due to the larger access granularity of SSDs, prefetching is more likely to result in cache pollution. Our prefetching decisions are based on the history of cache misses, with prefetching only executed when there is a high confidence level that it will yield benefits. Furthermore, given the higher access latency of SSDs, we employ the Best-Offset prefetching algorithm [44], which prioritizes timeliness in retrieving data from the backend SSD. When a memory access request reaches the CXL-SSD, it is directly returned if the SSD cache hits; otherwise, it is translated into SimpleSSD data packets for backend retrieval.
@@
We ran the Viper test on the CXL-SSD to evaluate its performance. Unlike the previous tests, we utilized the AM mode in this case to demonstrate its usage. As illustrated in Fig. 15, the test results indicate that the QPS of CXL-SSD across four operations are considerably lower than CXL-DRAM, primarily due to the higher access latency inherent in the Flash medium. To evaluate the effects of cache on system performance, we implemented and compared different cache replacement policies including least recently used (LRU) and first in first out (FIFO). It can be seen that the QPS of LRU or FIFO CXL-SSD with an additional cache has been significantly enhanced, reaching 72%-88% of CXL-DRAM. These results not only prove the effectiveness of the AM mode, but also demonstrate the high expandability of our simulator.
Nevertheless, the design of CXL-SSD as a part of memory pool still faces many challenges and trade-offs that warrant further investigation: (1) Bridging the performance gap between SSD and DRAM. The inherent high latency of flash media can severely degrade system performance, thus innovations are required to boost CXL-SSD performance (e.g., device-side cache, dedicated prefetch strategies). Moreover, designing an efficient tiered memory system is also pivotal for the practical deployment of CXL-SSD. (2) Coordination of CXL.io and CXL.mem transfers. Although CXL.mem provides lower-latency, memory-semantic access to the SSD, the DMA-based block transfer means supported by CXL.io must not be abandoned since it is superior in transferring data blocks larger than [45]. Efficiently coordinating both transfer channels at the software level remains a significant challenge. (3) Volatile versus non-volatile usage modes. CXL-SSD should not be simply used as a type of cheap and large-volume memory complementing the traditional volatile main memory, its persistent property can make a big difference. CXL-SSD and more broadly CXL-NVM bring a revival of persistent memory, which partly replaces the storage layer thus is vital to applications such as in-memory databases, big data analytics, and virtualization.
| Attribute | CXL-DMSim | QEMU | Mess+gem5 | gem5-cxl | DisaggSim | DRackSim | CXLMemSim | Remote NUMA |
| Reference | This work | [12] | [13] | [14] | [15] | [16] | [17] | [18, 19] |
| CXL protocol support | 1.1+ | 2.0 | No | No | No | No | No | No |
| Full-system CXL support | Yes | Yes | No | No | No | No | No | No |
| Silicon validation | Yes | No | No | No | No | No | No | N/A |
| Expandability | High | High | Low | Low | Medium | Medium | Low | Low |
| Sw/hw co-design | Yes | Yes | No | No | No | No | No | No |
| Cycle accuracy | Yes | No | Yes | Yes | Yes | Yes | No | Yes |
| Development maturity | Yes | Yes | Yes | No | Yes | Yes | No | Yes |
| Configurability | High | High | High | Low | High | High | Medium | Low |
| Simulation error | Low | High | Medium | High | High | High | Unknown | High |
| Simulation speed | Low | Medium | Low | Low | Low | Low | Medium | High |
@@ V-A Simulation Overhead of CXL-DMSim
To assess the simulation overhead covering two aspects which are simulation speed and occupied host memory introduced by CXL-DMSim in comparison to the native gem5, we conducted various experiments on our simulator and collected statistics of the execution time (HostSeconds) and physical memory occupation (HostPhyMemory) on the host machine. Fig. 16(a) compares the HostSeconds between native gem5, CXL-DMSimL, and CXL-DMSimA with respect to three workloads. The execution time of CXL-DMSimL is nearly identical to that of native gem5, while the average execution time of CXL-DMSimA is 16% higher. This is because CXL-DMSimA involves more components and longer latency in the simulation process. Notably, LMbench, a benchmark focused on evaluating memory access latency, exhibits slightly higher execution times compared to STREAM and MERCI.
Fig. 16(b) compares HostPhyMemory overhead. CXL-DMSim exhibits an average increase of only 3.5% in host memory usage, compared to raw gem5. This increase results from the additional memory allocation necessary on the host machine to model the CXL device memory. Moreover, the magnitude of memory overhead is closely correlated with the amount of physical memory utilized by different applications.
@@ V-B Simulation Error of CXL-DMSim
We also analyzed the simulation error stats of CXL-DMSim for a large variety of workloads against the results on our hardware testbed. The collected error data is compared with that of other CXL simulators as depicted in Fig. 17. It can be seen that the minimum simulation error is 1.9% for MERCI, while the maximum goes to Viper. The averaged simulation error is measured at 3.4% across all workloads in our experiments. CXL-DMSim exhibits the best accuracy in comparison to the two latest works Mess (6.0%) and gem5Tune (6.2%) as revealed in their respective papers [13, 34]. The native gem5 shows the highest average error of 15.0% [13].
The sources of simulation errors for CXL-DMSim can be attributed to the cumulative effects of modeling inaccuracies in the entire system. These include but are not limited to the following. (1) The CPU model lags behind the latest CPU we used on our testbed. (2) The cache model accounts for a large part of errors, particularly when data volumes range from one to four times the LLC capacity. (3) The poor modeling of the disk in gem5 when storage layers are involved, as can be observed with the increased errors for Viper when some data was pushed out to disk. Note that it is always a dilemma in the pursuit of simulation accuracy and speed.
@@
Table V compares CXL-DMSim with other prominent simulators at various aspects. First, CXL-DMSim and QEMU offer comprehensive support for the CXL protocol, which is critical for accurately simulating the behavior of real CXL memory systems. In contrast, the other six simulators actually do not support the CXL protocol. Notably, although DisaggSim and DRackSim support multi-node interconnection simulation, they lack of modeling the CXL protocol’s specific workflow and device behavior. Second, CXL-DMSim is a full-system CXL simulator that provides the most realistic interaction between the operating system and simulated hardware, opening up a range of possibilities for software-hardware co-design. In contrast, Mess, gem5-CXL, CXLMemSim, DisaggSim, and DRackSim cannot boot the operating system that supports CXL devices. Third, CXL-DMSim is the only one which has undergone extensive silicon validation, surpassing the other simulators in terms of accuracy and realism. CXL-DMSim also excels in expandability, cycle accuracy, development maturity, and configurability with acceptable simulation error. Nevertheless, these advantages come along at the cost of slow simulation speed, high modeling complexities for new features, and limited system scale.